OpenStudy (anonymous):

30. Find the vector v with the given magnitude and the same direction as u. ||v|| = 3 u = (4,-4)

OpenStudy (anonymous):

help?

jimthompson5910 (jim_thompson5910):

Hint: Find the unit vector pointing in the same direction as u = (4, -4) and scale that unit vector by 3 to get your answer.

OpenStudy (anonymous):

In the same direction means just find a unit vector in the direction of 4,-4 then multiply it by the magnitude, 3. So to do that divide by the magnitude: \[||u||=\sqrt{16+16}=\sqrt{32}\] Means: \[\vec{w}_{\vec{u}}=\left< \frac{4}{\sqrt{32}},\frac{-4}{\sqrt{32}} \right>\] So: \[\vec{v}=3*\vec{w}_{\vec{u}}=\left< \frac{12}{\sqrt{32}},\frac{-12}{\sqrt{32}} \right>\]

OpenStudy (anonymous):

is that all that i would put in my answer??

OpenStudy (anonymous):

That's all I would. I don't know if you need to explain your procedure.

OpenStudy (anonymous):

Also note that: \[\vec{w}_{\vec{u}}\] Denotes the unit vector in the u direction.

OpenStudy (anonymous):

ok somehow i got the answer as \[\left(\begin{matrix}3\sqrt{2} \\8\end{matrix}\right)\]

OpenStudy (anonymous):

i keep on thinking that its wrong...

OpenStudy (anonymous):

But I just worked it all out above...check your work against mine...

OpenStudy (anonymous):

ohh ok i see i didnt't finish my work. i had to write out v = part does that have to be in that fraction form or decimal form?

OpenStudy (anonymous):

thanks!!!!!!!!!! :D

OpenStudy (anonymous):

I would do fraction.